### v1

# AI-based particle track identification in scintillating fibres read out with imaging sensors

Abstract
This paper presents the development and application of an AI-based method for particle track identification using scintillating fibres read out with imaging sensors. We propose a variational autoencoder (VAE) to efficiently filter and identify frames containing signal from the substantial data generated by SPAD array sensors. Our VAE model, trained on purely background frames, demonstrated a high capability to distinguish frames containing particle tracks from background noise. The performance of the VAE-based anomaly detection was validated with experimental data, demonstrating the method’s ability to efficiently identify relevant events with rapid processing time, suggesting a solid prospect for deployment as a fast inference tool on hardware for real-time anomaly detection. This work highlights the potential of combining advanced sensor technology with machine learning techniques to enhance particle detection and tracking.
1 Introduction
Scintillating fibre (SciFi) detectors are among the innovative concepts currently being explored for future particle detection. Offering excellent spatial resolution, these detectors are critical for enhancing the reconstruction of low-momentum particles [1]. However, constructing large detector volumes with SciFi would lead to an incredibly high number of channels if read out with silicon photomultipliers (SiPMs). Hence, different SciFi readout schemes based on imaging sensors have been used. For example, in Ref. [2], CCD cameras were coupled with image intensifier tubes. Recent work demonstrated particle tracking with SciFi read out by a CMOS single-photon avalanche diode (SPAD) array sensor [3], illustrating the feasibility of reading out a SciFi volume with single fibre resolution and minimising the number of electronics channels. However, the significant data volume and the challenge of distinguishing frames containing signal from frames containing background (i.e., dark counts or crosstalk) remain issues. As a form of anomaly detection, machine learning can be used to learn from background samples alone, identifying signal events based on their deviation from the background. This approach is more robust than traditional methods that require a predefined mix of background and signal samples for training, and it is typically much faster compared to traditional solutions [4]. In this paper, we present an AI-based method using a variational autoencoder (VAE) to identify particle tracks in scintillating fibres read out with a SPAD array sensor. The method has been tested on the data acquired in Ref. [3]. The VAE is trained exclusively on background frames to learn the typical noise pattern, enabling it to identify frames containing signal events with higher accuracy. We evaluate the performance of our model using experimental data and demonstrate its ability to distinguish between signal and background frames effectively.
The rest of the article is structured as follows: Section 2 provides a brief overview of autoencoders and their application in anomaly detection. Section 3 presents the VAE architecture, training process, and methodology for anomaly detection; it also discusses the results of applying the VAE to experimental data, including evaluating its computational performance. Finally, Section 4 concludes the paper and outlines potential future work.
2 Autoencoders
An autoencoder is a self-supervised neural network that is trained to reconstruct an output which is close to its original input. It consists of two parts, an encoder and a decoder. The encoder $f$ transforms the input $\bm{x}$ into a latent representation $\bm{z}=f(\bm{x})$ and the decoder maps the latent representation $\bm{z}$ onto the reconstructed output $\bm{\hat{x}}=g(\bm{z})$. The difference between the input $\bm{x}$ and the reconstruction $\bm{\hat{x}}$ is called reconstruction error. The autoencoder learns to minimise this error. Thereby, the main goal is not for the autoencoder to reconstruct the input perfectly, but that it learns a useful representation of the data. This can be achieved by introducing some form of constraint, for example by forcing $\bm{z}$ to have lower dimension than $\bm{x}$ [5, 6, 7].
The concept of autoencoders has been generalised from deterministic encoder and decoder functions to stochastic models which learn encoder and decoder distributions. These models are called variational autoencoders [6].
2.1 Variational Autoencoders
Just like the standard autoencoder, a variational autoencoder (VAE) is composed of an encoder and a decoder which are trained to minimise a reconstruction error. However, as opposed to the standard autoencoder where the input is encoded as a single point, the variational autoencoder encodes the input as a distribution over the latent space, allowing the model to capture uncertainty in the data and creating a continuous and smooth latent space, among other benefits.
In more detail, the latent variable $\bm{z}$ of some input $\bm{x}$ is constrained to be a random variable which is distributed according to some prior distribution $p_{\bm{\theta}}(\bm{z})$. The true posterior distribution $p_{\theta}(\bm{z}|\bm{x})$ is intractable but it can be approximated by $q_{\bm{\phi}}(\bm{z}|\bm{x})$ using variational inference techniques. This sets up a family of parameterised distributions as priors, usually Gaussian distributions with mean 0 and unitary variance, and chooses the distribution with the smallest approximation error. The parameters are optimised according to
$\mathcal{L}(\bm{\theta},\bm{\phi},\bm{x})=\mathbb{E}_{q_{\Phi}(\bm{z}|\bm{x})}[p_{\theta}(\bm{x}|\bm{z})]-\mathcal{D}_{KL}(q_{\Phi}(\bm{z}|\bm{x})||p_{\theta}(\bm{z})).$ (2.1)
The first term is equivalent to the reconstruction error and measures the reconstruction quality. The second term is the Kullback-Leibler Divergence. This ensures that the approximation $q_{\bm{\phi}}(\bm{z}|\bm{x})$ is close to the true posterior distribution $p_{\theta}(\bm{z}|\bm{x})$. [8, 9].
Figure 1: Variational Autoencoder simplified structure.
In practice, this means, given an input $\bm{x}$, the network returns the parameters defining the probability distribution $q_{\bm{\phi}}(\bm{z}|\bm{x})$ over the latent space that is most likely to produce $\bm{x}$. Since this distribution was constrained to be Gaussian, the defining parameters are the mean $\bm{\mu_{\phi}}(\bm{x})$ and the variance $\bm{\sigma_{\phi}^{2}}(\bm{x})$ [9]. The latent variable $\bm{z}$ is then sampled from $q_{\bm{\phi}}(\bm{z}|\bm{x})$. However, backpropagation cannot deal with this random sampling. Therefore, the sampling is done via a reparameterisation trick by describing $\bm{z}$ as a differentiable function of another random variable $\varepsilon\sim\mathcal{N}(0,\mathrm{I})$. Then the latent variable is calculated as $\bm{z}=\bm{\mu_{\phi}}(\bm{x})+\bm{\sigma_{\phi}}(\bm{x})\odot\bm{\varepsilon}$ where $\odot$ is the Hadamard product [8]. The decoder reconstructs the input using the sampled $\bm{z}$ as for a standard autoencoder. This workflow is illustrated in Fig. 1.
2.2 Anomaly Detection with VAE
VAEs have shown success in various anomaly detection tasks [10, 9, 11, 12]. As discussed at the beginning of this section, autoencoders map an input into a lower dimensional latent space. Through this constraint, it is expected, that the autoencoder learns useful features of the data. It learns to reconstruct data frequently exposed during training but fails on rare instances in the training set. Thus, when training an autoencoder on a dataset, the reconstruction is expected to fail when encountering anomalous data [5].
However, the connection between anomalies and large reconstruction errors is not entirely clear [13]. Using VAEs has the advantage that the Kullback-Leibler Divergence term regularises the latent space by forcing it to follow a well-defined prior distribution. The model learns the distribution of latent variables instead of the variables itself. This means, we can label an event as anomalous if the probability of it being associated with the learned latent distribution is low. Therefore, instead of the reconstruction error, also the regularised latent space can be used for anomaly detection [4].
2.3 Anomaly Detection in Particle Physics
In collider experiments, anomaly detection is typically used to monitor detectors and identify unexpected events, potentially indicating new physics beyond the Standard Model (BSM). Machine learning methods, including VAEs, enable model-agnostic anomaly detection, enhancing the discovery potential for BSM events [14, 15, 4].
Recent advancements have shown that VAEs can be particularly effective in high-dimensional data typical of particle physics experiments [16, 17]. By learning the underlying distribution of the Standard Model events, VAEs are hypothesised to be capable of distinguishing between common physical processes and rare anomalies that may signal new physics. For instance, studies have demonstrated the ability of VAEs to identify anomalies in the Large Hadron Collider (LHC) data, aiding in the search for novel particles or interactions [18, 19].
Future research may focus on enhancing the interpretability of the latent space in VAEs, facilitating a deeper understanding of the detected anomalies. Additionally, integrating VAEs with other machine learning approaches, such as supervised learning algorithms, could further improve anomaly detection capabilities. The development of more sophisticated models and training techniques will likely continue to expand the applicability of VAEs in particle physics and other scientific domains [4].
3 Anomaly Detection with VAE
This section details the VAE model architecture, training process, and its application for anomaly detection.
3.1 VAE Architecture and Training Details
The VAE used in this study comprises an encoder with six convolutional layers, each followed by LeakyReLU activations with a negative slope of 0.01. The convolutional layers progressively increase from 8 to 64 filters, using $4\times 4$ kernels with a stride of 2, effectively downsampling the input while extracting features. The encoder output is flattened and passed through fully connected layers to produce the mean ($\bm{\mu}$) and log-variance ($\log(\bm{\sigma}^{2})$) vectors for the 32-dimensional latent space. The decoder uses a symmetrical structure with six transposed convolutional layers and LeakyReLU activations to upsample the latent vector back to the original input dimensions, with a final sigmoid activation ensuring output values in the range [0, 1]. The model contains 605,513 trainable parameters, balancing complexity with computational efficiency.
The loss function $\mathcal{L}$ combines the Binary Cross-Entropy (BCE) loss and Kullback-Leibler Divergence (KLD).
The BCE loss is given by:
$\mathcal{L}_{\text{BCE}}(\bm{x},\bm{\hat{x}})=-\sum_{n}\left[x_{n}\cdot\log(\hat{x}_{n})+(1-x_{n})\cdot\log(1-\hat{x}_{n})\right]$ (3.1)
where $x_{n}$ represents the ground truth binary label for the $n$-th pixel, and $\hat{x}_{n}$ is the predicted probability that the label is $1$. Here, $\bm{x}$ and $\bm{\hat{x}}$ are vectors containing the ground truth labels and predicted probabilities for all pixels, respectively.
The KLD term is:
$\mathcal{L}_{\text{KLD}}(\bm{\mu},\bm{\sigma})=-\frac{1}{2}\sum_{i}\left[1+\log(\sigma_{i}^{2})-\sigma_{i}^{2}-\mu_{i}^{2}\right]$ (3.2)
where $\bm{\mu}=(\mu_{1},\mu_{2},\dots,\mu_{d})$ and $\bm{\sigma}=(\sigma_{1},\sigma_{2},\dots,\sigma_{d})$ are the mean and standard deviation vectors of the latent variable’s Gaussian distribution. The summation is over the latent space dimensions, with $d$ being the dimensionality of the latent space.
The overall loss function $\mathcal{L}$ is then a combination of both terms:
$\mathcal{L}(\bm{x},\bm{\hat{x}},\bm{\mu},\bm{\sigma})=\mathcal{L}_{\text{BCE}}(\bm{x},\bm{\hat{x}})+\beta\cdot\mathcal{L}_{\text{KLD}}(\bm{\mu},\bm{\sigma})$ (3.3)
where $\beta$ is a weighting factor applied to the KLD term. In this work, the $\beta$ term is introduced following the approach outlined by [20], where it is scheduled using a sigmoid function during a 5-cycle annealing schedule [21]. This scheduling mitigates the issue of KLD vanishing and facilitates the formation of a well-organised latent space.
The VAE was implemented using PyTorch 1.14.0 [22] with Python 3.8.10 and trained on an NVIDIA V100 GPU with a batch size of 256 for 50 epochs. The model parameters were optimised using the AdamW optimiser with a learning rate of $1\times 10^{-3}$. To ensure a smooth training process, we employed a learning rate schedule that consisted of a linear warm-up phase followed by cosine annealing. The warm-up phase was applied during the first epoch, after which a cosine annealing schedule was used for the remaining epochs. This combined scheduling approach facilitated stable initial convergence while enabling efficient exploration of the loss landscape during later stages of training. The $\mathrm{BG}$ dataset consists of $5,012,736$ frames, split in the following way: $60\%$ for training, $10\%$ for validation, and $30\%$ for testing. The training and validation losses are shown in Fig. 2.
Figure 2: The plot illustrates the training and validation losses of a Variational Autoencoder (VAE) across iterations, with Binary Cross-Entropy (BCE) loss shown in red and Kullback-Leibler Divergence (KLD) loss in green. Solid lines represent validation losses, while dashed lines denote training losses. The learning rate (blue dotted line) and the $\beta$ weight of the KLD term (purple dash-dot line) are also plotted in arbitrary units.
3.2 Anomaly Detection
Although the VAE is trained on the entire BG dataset, the anomaly detection analysis specifically targets frames with four to ten counts. This range is chosen to maximise effectiveness because it represents the photon count where distinguishing the presence of a signal remains uncertain. Visually detecting signal tracks is extremely challenging due to the low photon counts per frame, making it difficult to distinguish them from background events that only contain DCs. However, by examining the total sum of frames for both the $\mathrm{BG}$ test set and the ${}^{90}\mathrm{Sr+BG}$ dataset (Fig. 3), we can slightly identify electrons traversing the scintillating material (from top to bottom in Fig. 3(b)). In contrast, the image in the BG case appears as pure noise (Fig. 3(a)).
(a) $\mathrm{BG}$ (test set).
(b) ${}^{90}\mathrm{Sr+BG}$.
Figure 3: Total frame sum within the fibre bundle region. Data taken from Ref. [3]
To identify such signal frames in the ${}^{90}\mathrm{Sr+BG}$ dataset, we utilise the loss values from our VAE. Figure 4 illustrates the distribution of BCE and KLD loss values for the test set of the $\mathrm{BG}$ and the ${}^{90}\mathrm{Sr+BG}$ dataset. The plots indicate the divergence points, where the loss distributions of the ${}^{90}\mathrm{Sr+BG}$ data start to exceed those of the BG test set. Both these divergence points and the 98th percentile of the $\mathrm{BG}$ test set, which is also marked in the plots, are used to identify anomalies.
(a) BCE loss.
(b) KLD loss.
Figure 4: Comparison of loss distributions for the Variational Autoencoder (VAE) model. The plot on the left shows the distribution of the Binary Cross-Entropy (BCE) reconstruction loss, while the right shows the distribution of the Kullback-Leibler Divergence (KLD) loss. In both plots, the blue curve represents the loss distribution for the $\mathrm{BG}$ test set, and the red curve represents the loss distribution for the ${}^{90}\mathrm{Sr+BG}$ data. The vertical dashed green lines indicate the divergence points, where the loss distributions of the ${}^{90}\mathrm{Sr+BG}$ data begin to exceed those of the $\mathrm{BG}$ test set. The vertical dashed-dotted purple lines mark the 98th percentile of the loss distribution for the $\mathrm{BG}$ test set, serving as a threshold to identify outliers or regions of interest where the loss values are notably high.
An example of our anomaly detection method is presented in Fig. 5, where we utilise the 98th percentile of the BCE and KLD losses from the $\mathrm{BG}$ test set to select frames in both the $\mathrm{BG}$ test set and the ${}^{90}\mathrm{Sr+BG}$ dataset. For the $\mathrm{BG}$ sample, the separation appears completely random to the naked eye (Fig. 5(a)). On the other hand, for the ${}^{90}\mathrm{Sr+BG}$, the figure illustrates the VAE’s ability to extract signal frames, where the pattern of short electron tracks traversing from top to bottom is distinctly visible, in contrast to the remaining candidates, where the summed frames resemble random noise (Fig. 5(b)).
(a) Anomaly selection on the test set of the $\mathrm{BG}$ sample.
(b) Anomaly selection on the ${}^{90}\mathrm{Sr+BG}$ sample.
Figure 5: Summed frames within the fibre bundle region, separated based on whether they are considered anomalies (either BCE or KLD loss exceeded the 98th percentile on the $\mathrm{BG}$ test set) or not.
To facilitate the detection of anomalies within the 32-dimensional latent space of the VAE, we employed two widely used dimensionality reduction techniques, Uniform Manifold Approximation and Projection (UMAP) [23] and t-distributed Stochastic Neighbor Embedding (t-SNE) [24]. These methods enable the projection of high-dimensional data into lower-dimensional spaces (2D in our case), allowing for the visualisation of latent space structures and identifying potential outliers (Fig. 6); in both plots, most of the points are organised in a circular shape—likely due to the Gaussianity enforcement in the latent representation—reflecting the VAE’s 32-dimensional latent space structure. The anomaly points are distinctly positioned along the borders of the the circles, deviating from the dense cluster of normal points. This spatial separation suggests that the latent space contains enough information to detect signal in the frames.
(a) UMAP.
(b) t-SNE.
Figure 6: 2D projections of the 32-dimensional latent space for the $\mathrm{BG}$ test set and the ${}^{90}\mathrm{Sr+BG}$ dataset. The anomaly candidates are tagged following the selection from Fig. 5.
Table 1 shows the correlation between the number of counts and selected frames. The percentage of selected frames as anomaly candidates is higher in the ${}^{90}\mathrm{Sr+BG}$ sample compared to the $\mathrm{BG}$ test set for every number of counts per frame. In most cases, it is significantly higher or at least the same. We can conclude that the selection of a frame is not due to its higher number of counts but rather due to the signal event it likely contains. Of course, depending on the goal of the analysis, one can lower the threshold used in the BCE and KLD losses to select anomalies to avoid missing signals, but at the cost of allowing more background frames (high selection efficiency) or minimise the background at the risk of missing some signals (high selection purity). Additional results using different configurations of loss terms and thresholds are provided in Appendix A. Both the KLD and BCE loss terms prove useful for anomaly detection.
Counts BG (test) ${}^{90}\mathrm{Sr+BG}$
Frames Selected % Frames Selected %
4 23396 104 0.44 153066 2944 1.92
5 4830 367 7.60 44146 13630 30.87
6 828 424 51.21 14397 12457 86.52
7 109 95 87.16 7521 7494 99.64
8 14 14 100.00 5768 5768 100.00
9 1 1 100.00 5159 5159 100.00
Table 1: Comparison of the fraction of selected frames for the $\mathrm{BG}$ test set and the ${}^{90}\mathrm{Sr+BG}$ sample across different numbers of counts per frame. Frames are identified as anomalies if either the BCE or KLD loss exceeds the 98th percentile on the $\mathrm{BG}$ test set.
3.3 Fast inference
The computational performance of the VAE for anomaly detection was evaluated on both CPU and GPU platforms, with the results summarised in Table 2. The computation times, reported in milliseconds per frame, illustrate the VAE’s efficiency in processing batches of sensor data. Both CPU and GPU platforms significantly reduce processing time per frame as batch size increases. However, the GPU consistently outperforms the CPU, achieving computation times up to two orders of magnitude lower, with the fastest time recorded at just 0.08 ms per frame for a batch size of 64. These findings underscore the VAE’s potential for real-time signal detection tasks, particularly when utilising GPU hardware, which provides substantial speedups over CPU processing.
Phase Device Batch size (number of frames processed in parallel)
1 2 4 8 16 32 64
Time per frame [ms]
Encoder CPU 3.69 $\pm$ 1.23 1.82 $\pm$ 0.58 1.28 $\pm$ 0.66 0.69 $\pm$ 0.38 0.35 $\pm$ 0.16 0.22 $\pm$ 0.11 0.10 $\pm$ 0.12
GPU 1.60 $\pm$ 7.50 0.72 $\pm$ 0.93 0.38 $\pm$ 0.13 0.21 $\pm$ 0.15 0.10 $\pm$ 0.05 0.05 $\pm$ 0.05 0.01 $\pm$ 0.02
Decoder CPU 19.05 $\pm$ 1.86 15.73 $\pm$ 0.72 15.42 $\pm$ 0.70 15.35 $\pm$ 0.46 15.08 $\pm$ 0.45 13.94 $\pm$ 4.34 6.94 $\pm$ 7.58
GPU 0.46 $\pm$ 6.96 0.31 $\pm$ 0.89 0.13 $\pm$ 0.11 0.09 $\pm$ 0.14 0.09 $\pm$ 0.05 0.08 $\pm$ 0.05 0.04 $\pm$ 0.05
KLD CPU 0.17 $\pm$ 0.02 0.09 $\pm$ 0.01 0.05 $\pm$ 0.01 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00
GPU 0.25 $\pm$ 1.43 0.10 $\pm$ 0.01 0.05 $\pm$ 0.01 0.03 $\pm$ 0.00 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00
BCE CPU 0.90 $\pm$ 0.29 0.54 $\pm$ 0.16 0.30 $\pm$ 0.06 0.16 $\pm$ 0.01 0.09 $\pm$ 0.01 0.07 $\pm$ 0.02 0.03 $\pm$ 0.03
GPU 0.13 $\pm$ 0.49 0.06 $\pm$ 0.01 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Memory allocation GPU 0.16 $\pm$ 0.04 0.11 $\pm$ 0.02 0.09 $\pm$ 0.01 0.07 $\pm$ 0.01 0.05 $\pm$ 0.00 0.04 $\pm$ 0.01 0.02 $\pm$ 0.02
Total CPU 23.81 $\pm$ 1.51 18.18 $\pm$ 0.74 17.05 $\pm$ 0.58 16.24 $\pm$ 0.57 15.53 $\pm$ 0.48 14.23 $\pm$ 4.43 7.07 $\pm$ 7.72
GPU 2.60 $\pm$ 2.50 1.31 $\pm$ 0.08 0.69 $\pm$ 0.05 0.41 $\pm$ 0.03 0.26 $\pm$ 0.01 0.18 $\pm$ 0.06 0.08 $\pm$ 0.08
Table 2: Comparison of computation times for different batch sizes on CPU (Intel(R) Xeon(R) CPU E5-2698 v4 @ 2.20GHz) and GPU (Tesla V100-SXM2) systems. The time reported is the VAE processing time per frame, and each result is an average of 1000 independent computations. The system has a total memory of 528 GB.
4 Conclusion
Scintillating fibres read out with imaging sensors offer an optimal configuration for high spatial resolution particle tracking. Given the growing success of machine learning techniques for anomaly detection, we proposed using a variational autoencoder (VAE) to differentiate between signal and background frames. In this study, we demonstrated that a VAE, trained solely on background frames, can serve as an effective and fast selection algorithm. We used the Binary Cross-Entropy (BCE) and the Kullback-Leibler Divergence (KLD) losses as anomaly scores to identify signal frames. The results showed that the frames selected indeed contained signal events. Furthermore, we tested the same method on background frames and found that significantly fewer frames were selected, confirming the effectiveness of the VAE with both BCE and KLD losses as anomaly detection metrics. Although we have not yet tested this approach with finer fibres, we hypothesise that the VAE would further improve its performance, given the lower number of dark counts and the more refined pattern left by a particle, which would be easier and faster for the method to detect. In future work, we plan to explore its integration as a real-time anomaly detection method into the hardware.
Appendix A Additional results
This appendix provides supplementary results, demonstrating the impact of varying thresholds on Binary Cross-Entropy (BCE) and Kullback-Leibler Divergence (KLD) loss values of the variational autoencoder (VAE). The results highlight how using these loss functions individually or in combination yields different sets of anomalies. Specifically, the following sections focus on anomaly selection based on the criteria detailed below:
  • Section A.1: Selection of anomalies using thresholds applied solely to the BCE loss.
  • Section A.2: Selection of anomalies using thresholds applied solely to the KLD loss.
  • Section A.3: Selection of anomalies using thresholds applied to both the BCE and KLD losses.
In each section, the following thresholds are applied to the respective loss functions to identify anomaly candidates, as discussed in Section  3:
  • Divergence threshold: The point at which the ${}^{90}\mathrm{Sr+BG}$ loss distributions start to exceed those of the $\mathrm{BG}$ test set, as illustrated in Fig. 4.
  • 98th percentile threshold: The 98th percentile of the loss distribution for the $\mathrm{BG}$ test set.
  • Max threshold: The maximum loss value observed in the $\mathrm{BG}$ test set.
A.1 Anomaly detection on BCE
Figure 8 shows the total sum of frames selected and not selected as anomaly candidates based on different thresholds applied to the BCE loss values, while Table 3 presents the total number of frames selected in each case, categorised by the number of photon counts in the frame. This example clearly illustrates that applying a higher threshold to the BCE loss value results in fewer background frames being selected as signal candidates, while more genuine signal candidates are overlooked; and vice versa.
Anomaly candidate frames summed.
Non-anomaly candidate frames summed.
(a) Divergence threshold (${}^{90}\mathrm{Sr+BG}$ > $\mathrm{BG}$ test set).
Anomaly candidate frames summed.
Non-anomaly candidate frames summed.
(b) 98th percentile threshold ($\mathrm{BG}$ test set).
Anomaly candidate frames summed.
Non-anomaly candidate frames summed.
(c) Max threshold ($\mathrm{BG}$ test set).
Figure 8: Summed frames within the fibre bundle region for the ${}^{90}\mathrm{Sr+BG}$ dataset, separated based on whether they are considered anomalies using the BCE loss.
Divergence threshold (${}^{90}\mathrm{Sr+BG}$ > $\mathrm{BG}$ test set)
Counts BG (test) ${}^{90}\mathrm{Sr+BG}$
Frames Selected % Frames Selected %
4 23396 1243 5.31 153066 16408 10.72
5 4830 1178 24.39 44146 18459 41.81
6 828 520 62.80 14397 12134 84.28
7 109 100 91.74 7521 7462 99.22
8 14 14 100.00 5768 5766 99.97
9 1 1 100.00 5159 5159 100.00
98th Percentile threshold ($\mathrm{BG}$ test set)
Counts BG (test) ${}^{90}\mathrm{Sr+BG}$
Frames Selected % Frames Selected %
4 23396 100 0.43 153066 2632 1.72
5 4830 207 4.29 44146 6828 15.47
6 828 197 23.79 14397 8448 58.68
7 109 66 60.55 7521 7114 94.59
8 14 13 92.86 5768 5762 99.90
9 1 1 100.00 5159 5159 100.00
Max threshold ($\mathrm{BG}$ test set)
Counts BG (test) ${}^{90}\mathrm{Sr+BG}$
Frames Selected % Frames Selected %
4 23396 0 0.00 153066 4 0.00
5 4830 0 0.00 44146 8 0.02
6 828 0 0.00 14397 177 1.23
7 109 0 0.00 7521 1350 17.95
8 14 0 0.00 5768 3490 60.51
9 1 0 0.00 5159 4750 92.07
Table 3: Comparison of the fraction of selected frames using the BCE loss for the $\mathrm{BG}$ testing set and the ${}^{90}\mathrm{Sr+BG}$ sample for different numbers of counts per frame.
A.2 Anomaly detection on KLD
Figure 10 shows the total sum of frames selected and not selected as anomaly candidates based on different thresholds applied to the KLD loss values, while Table 4 presents the total number of frames selected in each case, categorised by the number of photon counts in the frame. The visual separation between signal and background frames suggests that the KLD loss may not be as effective as the BCE loss for this particular case study. However, as shown in Fig. 6, the latent space clearly encodes signal events differently, indicating that further investigation would be required. This would be especially important if one aims to reduce the computational complexity of the VAE by running only the encoder and detecting anomalies based solely on the latent representation.
Anomaly candidate frames summed.
Non-anomaly candidate frames summed.
(a) Divergence threshold (${}^{90}\mathrm{Sr+BG}$ > $\mathrm{BG}$ test set).
Anomaly candidate frames summed.
Non-anomaly candidate frames summed.
(b) 98th Percentile threshold ($\mathrm{BG}$ test set).
Anomaly candidate frames summed.
Non-anomaly candidate frames summed.
(c) Max threshold ($\mathrm{BG}$ test set).
Figure 10: Summed frames within the fibre bundle region for the ${}^{90}\mathrm{Sr+BG}$ dataset, separated based on whether they are considered anomalies using the KLD loss.
Divergence threshold (${}^{90}\mathrm{Sr+BG}$ > $\mathrm{BG}$ test set)
Counts BG (test) ${}^{90}\mathrm{Sr+BG}$
Frames Selected % Frames Selected %
4 23396 1059 4.53 153066 33327 21.77
5 4830 2606 53.95 44146 31695 71.80
6 828 754 91.06 14397 12324 85.60
7 109 106 97.25 7521 6868 91.32
8 14 14 100.00 5768 5529 95.86
9 1 1 100.00 5159 5086 98.58
98th Percentile threshold ($\mathrm{BG}$ test set)
Counts BG (test) ${}^{90}\mathrm{Sr+BG}$
Frames Selected % Frames Selected %
4 23396 4 0.02 153066 321 0.21
5 4830 174 3.60 44146 7604 17.22
6 828 309 37.32 14397 7691 53.42
7 109 82 75.23 7521 5155 68.54
8 14 14 100.00 5768 4721 81.85
9 1 1 100.00 5159 4746 91.99
Max threshold ($\mathrm{BG}$ test set)
Counts BG (test) ${}^{90}\mathrm{Sr+BG}$
Frames Selected % Frames Selected %
4 23396 0 0.00 153066 0 0.00
5 4830 0 0.00 44146 0 0.00
6 828 0 0.00 14397 0 0.00
7 109 0 0.00 7521 18 0.24
8 14 0 0.00 5768 239 4.14
9 1 0 0.00 5159 893 17.31
Table 4: Comparison of the fraction of selected frames using the KLD loss for the $\mathrm{BG}$ testing set and the ${}^{90}\mathrm{Sr+BG}$ sample for different numbers of counts per frame.
A.3 Anomaly on both BCE and KLD
Figure 12 shows the total sum of frames selected and not selected as anomaly candidates based on different thresholds applied to either the BCE or KLD loss values, while Table 5 presents the total number of frames selected in each case, categorized by the number of photon counts in the frame. The results suggest that combining both BCE and KLD performs similarly to using BCE alone. Future work could explore more sophisticated methods of combining these loss functions to leverage their strengths and improve anomaly detection efficiency.
Anomaly candidate frames summed.
Non-anomaly candidate frames summed.
(a) Divergence threshold (${}^{90}\mathrm{Sr+BG}$ > $\mathrm{BG}$ test set).
Anomaly candidate frames summed.
Non-anomaly candidate frames summed.
(b) 98th Percentile threshold ($\mathrm{BG}$ test set).
Anomaly candidate frames summed.
Non-anomaly candidate frames summed.
(c) Max threshold ($\mathrm{BG}$ test set).
Figure 12: Summed frames within the fibre bundle region for the ${}^{90}\mathrm{Sr+BG}$ dataset, separated based on whether they are considered anomalies using both the BCE and the KLD losses.
Divergence threshold (${}^{90}\mathrm{Sr+BG}$ > $\mathrm{BG}$ test set)
Counts BG (test) ${}^{90}\mathrm{Sr+BG}$
Frames Selected % Frames Selected %
4 23396 2253 9.63 153066 47297 30.90
5 4830 3255 67.39 44146 39147 88.68
6 828 809 97.71 14397 14374 99.84
7 109 109 100.00 7521 7521 100.00
8 14 14 100.00 5768 5768 100.00
9 1 1 100.00 5159 5159 100.00
98th Percentile threshold ($\mathrm{BG}$ test set)
Counts BG (test) ${}^{90}\mathrm{Sr+BG}$
Frames Selected % Frames Selected %
4 23396 104 0.44 153066 2944 1.92
5 4830 367 7.60 44146 13630 30.87
6 828 424 51.21 14397 12457 86.52
7 109 95 87.16 7521 7494 99.64
8 14 14 100.00 5768 5768 100.00
9 1 1 100.00 5159 5159 100.00
Max threshold ($\mathrm{BG}$ test set)
Counts BG (test) ${}^{90}\mathrm{Sr+BG}$
Frames Selected % Frames Selected %
4 23396 0 0.00 153066 4 0.00
5 4830 0 0.00 44146 8 0.02
6 828 0 0.00 14397 177 1.23
7 109 0 0.00 7521 1365 18.15
8 14 0 0.00 5768 3620 62.76
9 1 0 0.00 5159 4833 93.68
Table 5: Comparison of the fraction of selected frames using both the BCE and KLD losses for the $\mathrm{BG}$ testing set and the ${}^{90}\mathrm{Sr+BG}$ sample for different numbers of counts per frame.
Acknowledgments
We would like to thank the AQUA-LAB at EPFL for their help and fruitful discussions about the data acquisition for this project. Further we would like to thank them for the provision of the SwissSPAD2 imaging sensor.
Part of this work was supported by the SNF grant PCEFP2_203261, Switzerland.